### v2

# Improving Representation of High-frequency Components for Medical Foundation Models

Abstract
Foundation models have recently attracted significant attention for their impressive generalizability across diverse downstream tasks. However, these models are demonstrated to exhibit great limitations in representing high-frequency components and fine-grained details. In many medical imaging tasks, the precise representation of such information is crucial due to the inherently intricate anatomical structures, sub-visual features, and complex boundaries involved. Consequently, the limited representation of prevalent foundation models can result in significant performance degradation or even failure in these tasks. To address these challenges, we propose a novel pretraining strategy, named Frequency-advanced Representation Autoencoder (Frepa). Through high-frequency masking and low-frequency perturbation combined with adversarial learning, Frepa encourages the encoder to effectively represent and preserve high-frequency components in the image embeddings. Additionally, we introduce an innovative histogram-equalized image masking strategy, extending the Masked Autoencoder (MAE) approach beyond ViT to other architectures such as Swin Transformer and convolutional networks. We develop Frepa across nine medical modalities and validate it on 32 downstream tasks for both 2D images and 3D volume data. Without fine-tuning, Frepa can outperform other self-supervised pretraining methods and, in some cases, even surpasses task-specific trained models. This improvement is particularly significant for tasks involving fine-grained details, such as achieving up to a +15 increase in dice score for retina vessel segmentation and a +7 increase in IoU for lung nodule detection. Further experiments quantitatively reveal that Frepa enables superior high-frequency representations and preservation in the embeddings, underscoring its potential for developing more generalized and universal medical image foundation models.
Index Terms:
Artificial intelligence, foundation model, autoencoder, segmentation, classification, detection.1 Introduction
Recently, the rapid development of medical foundation models has attracted considerable attention and emerged as a novel paradigm in artificial intelligence (AI). These models typically involve pretraining on unlabeled datasets using self-supervised learning techniques [1, 2, 3, 4] or employ paired images, annotations, and text for supervised training [5, 6]. By training on large-scale datasets, foundation models can exhibit emergent capabilities [7] and facilitate improved generalization by transferring the knowledge acquired during the pretraining phase to specific tasks or data sources. In the field of healthcare, many foundation models have achieved significant progress across various medical tasks [4, 5, 8, 6, 9], marking a substantial advancement in medical AI research.
However, both our findings and previous works suggest that these foundation models often result in limited representations of high-frequency components and fine-grained details [5, 10, 11]. Fig. 1 provides two illustrative examples. Fig. 1a shows the classification accuracy of different models on the MedMNIST dataset [12] after applying high-pass filters of varying sizes. The performance of these models degrades rapidly with increasing filter size, suggesting an over-reliance on low-frequency components and a limited ability to capture high-frequency components. Fig. 1b showcases the segmentation and detection results of tasks involving fine-grained details using the fine-tuned MedSAM [5] and our proposed methods. This highlights that even segmentation-specific foundation models struggle with fine-grained structures. In diagnostic practices, many diseases and lesions manifest as subtle differences from normal tissue or other disease subtypes, making accurate disease identification heavily reliant on detailed information confined to small regions [13]. Moreover, anatomical structures such as small lesions, blood vessels, and fibrotic tissues, which are critical for both physiological functions and disease diagnosis [14, 15], also involve high-frequency and fine-grained information. The limited representation of such information is an inherent shortcoming of current foundation models, leading to poor performance or even failure in these tasks and reducing their generalizability to real-world medical applications.
The limitations in high-frequency representation are primarily due to the dominant role of global information in current training strategies. For instance, autoencoder pretraining methods like the masked autoencoder (MAE) [1] typically produce reconstructions that capture only general semantic information while lacking specific details [11]. In fact, during the reconstruction process, low-frequency, global information tends to be more substantial, causing the model to emphasize these components over high-frequency details [16], resulting in limited detailed information. A similar issue is observed in contrastive learning-based methods [17]. Techniques such as CLIP [3] rely on textual information and label-level annotations [6], which contain more global-level information than pixel-level details, leading to a neglect of detailed information capturing. Additionally, as the commonly used backbone network, the vanilla vision transformer (ViT) exacerbates this issue. Compared to convolutional operators (e.g., ConViT [18]) and multi-scale feature maps (e.g., Swin Transformer, SwinT [19]), ViT has relatively weaker inductive biases, which results in a reduced ability to capture high-frequency components [20, 21]. This is also evident from Fig. 1a. Consequently, the inefficiency in capturing high-frequency information presents significant challenges for current foundation models in medical imaging tasks.
Motivated by the aforementioned observations, we introduce an innovative pretraining strategy for foundational models, termed the Frequency-advanced representation autoencoder (Frepa). Frepa incorporates a novel imaging corruption and reconstruction branch designed to improve the utilization and representation of high-frequency components. This branch, termed frequency dual-component masking, includes two key steps: 1) High-frequency domain masking: We randomly mask high-frequency domains in the frequency spectrum. During training, the image encoders use the remaining information to reconstruct the masked frequency components, encouraging the encoders to learn and leverage realistic high-frequency distributions. 2) Low-frequency perturbation: To prevent the image encoders from relying excessively on easily learned low-frequency components, we introduce a zero-mean perturbation to the low-frequency part of the spectrum, prompting the encoders to rely more heavily on the high-frequency components. Additionally, we propose a hierarchical frequency-to-spatial loss and design an adversarial training strategy to achieve a more robust image reconstruction. We also modify the MAE by incorporating histogram-equalized masking methods, extending its application beyond vanilla ViT to convolutional networks and multi-scale ViT architectures. Furthermore, we integrate the 2D pretrained encoder with iterative intra- and inter-image transformer blocks, enabling its direct application to 3D volume data without architectural modifications. Frepa is developed on 17 million images and validated on 32 downstream tasks across 9 modalities. Frepa not only demonstrates superior performance in general medical tasks but also shows more significant improvement in tasks involving fine-grained information, where other models yield almost failed results. For instance, Frepa outperforms the best other models by approximately 15 in dice similarity coefficient (DSC) for retina vessel segmentation, 12 in DSC for pulmonary artery-vein segmentation, and 7 in Intersection over Union (IoU) for pulmonary nodule detection. Additionally, we also design a quantitative experiment to directly demonstrate that Frepa enables better representation and preservation of high-frequency components in the image embedding compared to other autoencoder pretraining methods. These results showcase that Frepa outperforms in capturing high-frequency components and fine-grained details without compromising the representation of global-level semantic information. Our codes are publicly available as https://github.com/Arturia-Pendragon-Iris/Frepa.
Our contributions to the state-of-the-art (SOTA) are as follows:
We propose Frepa, a novel approach that significantly enhances the high-frequency capturing capability of foundation models without compromising their low-frequency representation.
We extend the MAE beyond vanilla ViT to improve its generalizability. Additionally, we extend pretrained 2D encoders for direct deployment on 3D volume data.
We develop Frepa on 17 million images and validated it on 32 downstream tasks for both 2D and 3D data, without requiring fine-tuning.
Extensive experiments demonstrate that Frepa outperforms SOTA methods on most downstream tasks, particularly those involving fine-grained information. Frepa also exhibits exceptional performance and generalizability on previously unseen modalities.
2 Related Work
2.1 Foundation model
The recent advancements in foundation models signify a transformative leap in deep learning, showcasing exceptional generalization and transfer learning capabilities. The multi-modal visual foundation models based on contrastive learning, such as CLIP [3], Flamingo [22], and LLaVA [23], have made significant strides in efficient transfer learning and exhibit impressive zero-shot performance. These models, however, rely on large-scale image-text paired datasets and are limited in their applicability to other visual tasks like segmentation or image restoration. On the other hand, autoencoder pretraining strategies enable the network to learn meaningful feature representations from large-scale data in a self-supervised manner via reconstructing raw images from corrupted images. Among these approaches, MAE [1] is particularly notable, which pretrains a ViT encoder by randomly masking parts of an image and reconstructing the original image from the remaining patches. However, recent studies have highlighted the limitation of autoencoders in that they tend to focus more on low-frequency information and fail to effectively utilize the full spectrum of frequency information [10, 11]. This issue is especially critical in medical imaging tasks, where accurate diagnosis and numerous physiological structures often depend on high-frequency components and fine-grained details. To the best of our knowledge, Frepa is the first study to systematically investigate the frequency representation capabilities of foundation models and to validate the feasibility in the medical field.
2.2 Medical foundation model
As foundation models for natural images have achieved significant success, their applications in the medical field are gaining increasing attention. Current research on medical foundation models can be broadly classified into specialist models and generalist models. Specialist models are typically designed to address clinical tasks within a single imaging modality or specific disease type. Examples include USFM [4] for ultrasound, RETFound [24] for retinal images, and EndoFM [25] for endoscopy videos. Although specialist models can achieve higher accuracy and efficiency within their specific modalities, they often face challenges in generalizing to other modalities. In contrast, generalist models, such as MedSAM [5], MedDr [6], and LLAVA [23], offer broader applicability and generalizability across various modalities and tasks. These methods rely on large amounts of training data, computational resources, and parameters, achieving progress in generalization. However, in the process of transferring pretraining techniques from natural images, these models often don’t adequately consider the data characteristics of medical images, such as fine anatomical structures, multi-scale objects, and complex, uncertain boundaries. Consequently, they may perform poorly on medical imaging tasks that depend on fine-grained details and high-frequency information [5, 6] (Fig. 1), posing a significant barrier to their practical application in clinical settings. In this work, we focus on lightweight generalist models and introduce a novel pretraining strategy designed to enhance the ability of foundation models to represent high-frequency information. Our approach aims to improve performance on fine-grained tasks, thereby extending the practical applicability and generalizability of these models in clinical practice.
3 Methods
3.1 Frepa pretraining strategy
3.1.1 Frequency dual-component masking
To enhance image understanding and high-frequency representation, we introduce the frequency dual-component masking strategy. This approach consists of two steps: high-frequency masking and low-frequency perturbation. By predicting image components of the masked frequencies, the image encoder can learn the high-frequency components and their realistic distributions. Simultaneously, the low-frequency perturbation prevents the encoder from overly relying on simple low-frequency components, instead encouraging it to utilize medium-frequency and high-frequency components for image reconstruction. This pretraining strategy can be seamlessly applied to both vanilla ViT, multi-scale ViTs, and convolutional networks without requiring changes to the framework (Fig. 2,a).
High-frequency domain masking Given an image , we can obtain its frequency spectrum via the 2D discrete centered Fourier transform:
Here, the subscript denotes the value of the image at the coordinate , and similarly, the subscript denotes the frequency spectrum at the coordinate . Masking the low-frequency components can result in significant image distortion that is extremely challenging for reconstruction; therefore, we adopt a progressive masking strategy. Our goal is to make patch masking more likely to occur in the high-frequency components and less likely in the low-frequency components. To achieve this, we set the size of patches in the frequency spectrum to and use an exponential decay function to control the masking probability (Fig. 2,b).
Here means the masking probability of patch , and is the Euclidean distance from the center of patch to the center of the frequency spectrum. is the cut-off distance, and we set (Fig. 2,b).
Low-frequency perturbation After the aforementioned stage, the high-frequency components are severely distorted. However, most low-frequency components remain unaffected because is zero at the center patch. To further reduce the dependence of the image encoder on the low-frequency components, we introduce low-frequency perturbation. Specifically, we add intensity-attenuated, zero-mean noise to the masked frequency spectrum:
Here is the Gaussian distribution, and is the standard deviation. We set to make the root mean square error (RMSE) of perturbed images maintaining approximately the same with other image masking pretraining methods [2, 26]. We also set the center of the noise to zero to ensure that the mean value of the image remains unchanged before and after the perturbation. Following these two stages, we perform the inverse Fourier transform on the corrupted frequency spectrum to obtain the degraded image.
3.1.2 Equaled-histogram image-domain masking
Although advanced multi-scale ViTs and convolutional networks have demonstrated superior high-frequency capturing abilities compared to the vanilla ViT [20, 19], the success of MAE relies on the global property and asymmetric structure of the vanilla ViT, which enables the application of the self-attention mechanism to discrete, non-overlapping image patches [27]. In contrast, multi-scale ViTs and convolutional networks typically incorporate local window operators, such as convolutional layers, which prevent them from this same capability. It has been shown that directly setting all masked patches to zero is inefficient [28], as this approach significantly distorts the image histogram distribution (Fig. 3). Additionally, the convolutional operator can result in the ”mask pattern vanishing” issue [28], causing inconsistent masking effects and ratios across different encoding layers.
To address this problem, we propose an innovative and straightforward image masking strategy, termed the histogram-equalized masking strategy. Specifically, we first uniformly sample 30 of the visible image patches, each with a size of . Unlike MAE [1], we do not discard the masked patches; instead, we replace them with random noise that has the same histogram distribution to the original patches. Denoting the one masked patch as , such an operation can be formulated as
where is the histogram-equalized noise, and denotes the histogram distribution. Practically, such distributions can be approximated by a Gaussian distribution with the same mean and variance as the original patches. Images corrupted via this strategy can keep the histogram consistency before and after the masking (Fig. 3). Notably, the frequency dual-component masking can also preserve the histogram distribution.
3.2 Pretraining pipeline of Frepa
Given a raw 2D image , we randomly use the frequency-domain mask and the image-domain masking with equal probability to perform corruption and obtain the degraded image . Our objective is to reconstruct from using an encoder-decoder network :
Following the general autoencoder pretraining approach [1, 2, 10], we utilize lightweight image decoders for the network, composed of only several convolutional layers. To better guide the image encoder in leveraging high-frequency components and prevent key high-frequency details from being diluted by low-frequency components, we employ a high-frequency augmentation strategy. Specifically, we manually extract the high-frequency and edge features from the raw image via a filter operator , and then concatenated it with the degraded image along the channel dimension. Then the pretraining process can be formulated as
To prevent leakage of high-frequency components, the filter used cannot be a frequency-domain filter, such as exponential filtering; instead, filters or edge detection functions performed in the image domain are more appropriate. In our experiments, we use the Hessian filter [29] to accomplish this task. For simplicity, we use instead of to represent the reconstructed image in the following text.
Loss function for reconstruction
We utilize the RMSE to guide the reconstruction in the image domain.
We also incorporate the image gradient loss to restrict the reconstruction of details and textures.
where is the gradient operator.
To further enhance the representation of high-frequency components, we aim for the reconstructed images to achieve satisfactory alignment with the raw images in the frequency domain. Initially, we employ the commonly used focal frequency loss (FFL) [30], specifically designed to enhance attention to high frequencies. However, in practice, we found that directly incorporating the FFL can result in aliasing artifacts in the reconstructed images (Fig. 4), which cannot be ideal reconstruction outcomes. We tend to consider that these artifacts may arise from the degeneration of high-frequency information due to overfitting in the frequency domains [31], as FFL could impose excessively strong constraints on high frequencies [4]. To address this issue, we propose a novel hierarchical frequency-to-spatial loss (HFL), which can be formulated as
denotes the high-pass exponential filter, while a higher represents less preservation of the low-frequency components. This loss function enables the optimization of high-frequency reconstruction in the image domain without directly imposing restrictions on the frequency spectrum, thereby preventing the formation of artifacts.
To further enhance the representation of high-frequency components, we propose an adversarial training strategy specifically designed for frequency dual-component masking. This strategy encourages the image embeddings to remain consistent under different masking and perturbation samplings. Specially, given two different masking and perturbation sampling, denoting as and , the corresponding degraded images are and , respectively. The objective function for the adversarial training is as follows:
where denotes the image encoders and is the Jensen–Shannon divergence. The total loss functions can thus be formulated as follows:
In our experiments, we set . For the images corrupted with the image-domain masking, is only calculated on the masked patches, which is consistent with MAE [1]; otherwise, the loss function is performed on the whole image. To enhance stability during the early stages of training, adversarial training is excluded for the first 5 epochs. Following this initial phase, the model is then trained using the previously described setting.
3.3 Extending to 3D volume decoding
The network and pretraining strategy above are designed for 2D images. However, in clinical practice, many datasets consist of 3D volumes and image sequences, such as computed tomography (CT) and magnetic resonance imaging (MRI) scans. The current image encoder architecture cannot be directly applied to these types of data. To address this challenge, inspired by the design of spatial-temporal networks [32], we have developed an image-fusion architecture to seamlessly transfer pretrained 2D encoders to 3D volume data (Fig. 2 c).
Given a volume data , where and represent the number of images and channels, respectively, we first apply the encoder to each image to obtain image embeddings with feature channels and embedding size . We then incorporate the intra-image position embedding into the embeddings as , and feed them into an inter-image transformer block to extract intra-image features, resulting in output , where is the token dimension. The transformer block follows the commonly used architecture [33]. is resized to and then processed by an inter-image transformer block, incorporating an inter-slice position embedding , to fuse the extracted features between contextual images. These two transformer blocks are iteratively applied to the embeddings to integrate intra- and inter-image information (with three iterations in our setting). The final output is then passed to a decoder head for various downstream tasks.
| Modality | Dataset | Metric | MAE | MFM | CLIP | MedSAM | MedDr | Frepa+ViT | Frepa+SwinT |
| CT | Chest cancer[34] | ACC | 0.888 | 0.794 | 0.889 | 0.857 | - | 0.937 | 0.920 |
| AUC | 0.983 | 0.964 | 0.982 | 0.973 | - | 0.991 | 0.983 | ||
| X-Ray | COVID19 Radiography[35] | ACC | 0.962 | 0.949 | 0.966 | 0.966 | - | 0.964 | 0.968 |
| AUC | 0.999 | 0.996 | 0.997 | 0.997 | - | 0.998 | 0.998 | ||
| Pneumonia MNIST[12] | ACC | 0.919 | 0.918 | 0.939 | 0.903 | 0.870 | 0.934 | 0.932 | |
| AUC | 0.973 | 0.966 | 0.982 | 0.965 | - | 0.970 | 0.981 | ||
| COVID19 PA Dataset[36] | ACC | 0.934 | 0.908 | 0.943 | 0.897 | - | 0.939 | 0.941 | |
| AUC | 0.980 | 0.980 | 0.981 | 0.977 | - | 0.984 | 0.984 | ||
| Lung cancer[37] | Detection | 0.911 | 0.908 | 0.843 | 0.899 | - | 0.951 | 0.948 | |
| IoU | 0.596 | 0.607 | 0.482 | 0.566 | - | 0.656 | 0.677 | ||
| OCT | OCT2017[38] | ACC | 0.815 | 0.866 | 0.868 | 0.837 | - | 0.888 | 0.880 |
| AUC | 0.932 | 0.960 | 0.972 | 0.944 | - | 0.970 | 0.963 | ||
| OCT MNIST[12] | ACC | 0.708 | 0.737 | 0.670 | 0.661 | 0.675 | 0.793 | 0.786 | |
| AUC | 0.903 | 0.913 | 0.908 | 0.880 | - | 0.940 | 0.932 | ||
| Ultrasound | Fetal Plane[39] | ACC | 0.928 | 0.922 | 0.932 | 0.919 | - | 0.897 | 0.919 |
| AUC | 0.959 | 0.959 | 0.946 | 0.955 | - | 0.945 | 0.951 | ||
| Breast MNIST[12] | ACC | 0.852 | 0.814 | 0.852 | 0.794 | 0.718 | 0.846 | 0.846 | |
| AUC | 0.861 | 0.834 | 0.833 | 0.823 | - | 0.863 | 0.798 | ||
| Retina | Aptos2019[40] | ACC | 0.764 | 0.796 | 0.758 | 0.788 | - | 0.802 | 0.808 |
| AUC | 0.882 | 0.908 | 0.891 | 0.893 | - | 0.914 | 0.906 | ||
| Dermatoscope | ISIC2019[41] | ACC | 0.701 | 0.696 | 0.705 | 0.701 | - | 0.715 | 0.717 |
| AUC | 0.887 | 0.902 | 0.892 | 0.891 | - | 0.921 | 0.916 |
| Modality | Dataset | Metric | MAE | MFM | CLIP | MedSAM | MedDr | Frepa+ViT | Frepa+SwinT |
| CT | COPD# | ACC | 0.869 | 0.828 | 0.878 | 0.801 | - | 0.887 | 0.882 |
| AUC | 0.946 | 0.933 | 0.930 | 0.929 | - | 0.954 | 0.960 | ||
| Pneumonia# | ACC | 0.867 | 0.842 | 0.897 | 0.882 | - | 0.897 | 0.887 | |
| AUC | 0.924 | 0.913 | 0.933 | 0.924 | - | 0.935 | 0.929 | ||
| Luna16 Lung Nodule[42] | Detection | 0.885 | 0.872 | 0.775 | 0.876 | - | 0.905 | 0.920 | |
| IoU | 0.569 | 0.548 | 0.410 | 0.518 | - | 0.626 | 0.633 |
4 Experiments and results
4.1 Implementation details
We conducted experiments on two popular architectures: vanilla ViT-B [43] (86.9M parameters) and SwinT-B [19] (86.9M parameters) as the image encoders. The image decoder for reconstruction consisted of four convolutional layers, while the classification head was composed of a multilayer perception. The models were pretrained on a comprehensive dataset, including modalities of CT, MRI, X-ray, ultrasound, optical coherence tomography (OCT, projection image), retina imaging, and dermatoscopy.111The sources of the datasets are provided on the GitHub webpage.. All images were resized to using bi-cubic interpolation and zero-padding. For single-channel grayscale images, the image was duplicated along the channel dimension to maintain consistency. The model was trained using the Adam optimizer with a learning rate of . We employed image augmentation techniques including random flipping, random rotation, random affine transformations, and random histogram shifting during the pretraining phase. The models were trained on a Linux workstation with two A100 (80G) GPUs for 100 epochs.
We compared Frepa with other pretrained and task-specific foundation models, including MAE [1], MFM [2] (a self-supervised strategy performed in the frequency domain), MedSAM [5], and CLIP [3]. The backbones for MAE and MFM were ViT-B, consistent with Frepa, and were pretrained on our collected dataset. The training of CLIP followed the method proposed in [6]. For MedSAM, we used the model and checkpoint published in [5] without modification. We also included one large visual-language model, MedDr [6] (40 billion parameters) for the comparison of some classification tasks.
4.2 Experiment settings for downstream tasks
We froze the encoders and fine-tuned only the task heads for downstream tasks. To ensure a fair comparison, we employed consistent training strategies and decoder architectures across all foundation models. The loss functions used were Cross Entropy loss for classification, Dice loss for segmentation, and GIoU loss [44] for detection tasks. For classification evaluation, we utilized accuracy (ACC) and the area under the curve (AUC) metrics. Segmentation performance was assessed using the DSC. Detection ratio and IoU were used for the evaluation of detection tasks. Notably, we incorporated two modalities—OCT reconstruction data and electron microscopy (EM) data—for segmentation tasks. These modalities were not included in the pretraining phases, thereby evaluating the generalizability to real-world data. The raw datasets were randomly divided into training, validation, and testing sets in an 8:1:1 ratio. Other training settings remain the same as above.
4.3 Visualization for reconstruction
We present several visualizations of reconstruction results on external datasets. Specifically, we compare reconstruction for image masking against MAE and reconstruction for frequency masking, with low-pass filtering, against MFM. As shown in Fig. 5, Frepa demonstrates the highest fidelity in reconstructing original images, with exceptional detail information. In contrast, images generated by other autoencoders appear relatively blurred, lacking sufficient detail in textures and edges. These results are further corroborated by the frequency spectrum and RMSE metrics of each image. The superior reconstruction quality and clear detail retention indicate that Frepa effectively captures medium- and high-frequency information.
4.4 Comparison with previous methods
4.4.1 Classification and detection
Tables I and II present the results of classification and lung nodule detection on 2D images and 3D volumes, respectively. It shows that Frepa, when implemented with both ViT and SwinT, overall outperforms MAE and MAF. Specifically, Frepa exhibits an average improvement on ACC of 2.3 and 2.4 over MAE, and 3.0 and 3.1 over MFM, respectively. AUC metric also shows a similar conclusion. Furthermore, when compared to the classification-specific foundation model, CLIP, Frepa demonstrates comparable performance across most tasks and excels in certain areas such as the classification of chest cancer sub-types and pneumonia sub-types. Additionally, it is noteworthy that in fine-grained tasks, such as lung nodule and lung nodule detection, Frepa achieves more substantial improvements, attributed to its superior capability of capturing high-frequency details. Specifically, For 2D images, there is a 4.0 increase in lung cancer detection ratio and a 7 improvement in IoU (Table I), while for 3D volumes, the lung nodule detection ratio improves by 3.5 and the IoU by 6.4 (Table II). Collectively, in classification tasks reliant on global-level information, Frepa achieves performance comparable to CLIP and MAE, affirming its robust global-level representation. Moreover, for tasks requiring fine-grained information, Frepa consistently outperforms, underscoring its effectiveness in capturing detailed, high-frequency information.
4.4.2 Segmentation
Tables III and IV present the segmentation results on 2D images and 3D volumes, respectively. These results encompass both large-scale anatomical segmentations, (lung, heart, and pneumonia infection in 2D segmentation, Tables III, and abdomen1K [45], AMOS22 [46], and ADCD [47] in 3D segmentation, Tables IV), as well as fine-grained and vascular-structure segmentation (OCT artery-vein [48], retinal vessels [49, 50, 51], and intercellular gaps [14] in 2D segmentation, Tables III, and pulmonary artery-vein [15] in 3D segmentation, Tables IV). In large-scale anatomical segmentations, Frepa shows a slight improvement over other models, including MedSAM, attributed to its enhanced ability to capture high-frequency details, thereby improving boundary identification. For the fine-grained segmentation tasks, Frepa exhibits more markedly superior performance, with a significant improvement ranging from +7 to +15 in DSC compared to the best-performing models. Notably, in vessel segmentation for retina and OCT (Tables III), other models yield almost failed segmentation results (Fig. 1), whereas Frepa consistently achieves exceptional results, maintaining topological structures and morphology that closely match the ground truth. Moreover, this advancement is preserved when applied to external modalities, such as the OCT reconstruction dataset and EM dataset, which exhibit entirely different imaging characteristics from other modalities. These findings robustly demonstrate Frepa’s superior representation of fine-grained information and high-frequency components, along with better generalizability than other foundation models when applied to a broader range of realistic medical tasks.
| Modality | Dataset | Metric | MAE | MFM | CLIP | MedSAM | Frepa+ViT | Frepa+SwinT |
| CT | Lung# | DSC() | 93.241.41 | 93.092.44 | 78.5314.62 | 95.363.47 | 96.891.19 | 96.462.80 |
| Heart# | DSC() | 88.682.02 | 88.972.30 | 65.1812.52 | 91.382.32 | 91.302.70 | 90.722.59 | |
| Pneumonia infection# | DSC() | 78.309.73 | 75.4710.18 | 55.1815.52 | 71.0710.72 | 81.478.19 | 81.519.35 | |
| OCT reconstruction | OCTA artery[48] | DSC() | 74.865.94 | 64.357.29 | 13.061.85 | 71.595.13 | 78.006.44 | 79.085.79 |
| OCT vein[48] | DSC() | 70.424.42 | 63.444.98 | 12.961.21 | 66.313.88 | 76.155.11 | 77.354.20 | |
| OCT artery[48] | DSC() | 53.708.39 | 46.458.67 | 13.101.59 | 54.517.50 | 59.456.41 | 61.886.95 | |
| OCT vein[48] | DSC() | 55.947.33 | 48.2510.40 | 13.090.96 | 53.308.11 | 60.757.36 | 61.668.52 | |
| Retina | HRF[49] | DSC() | 62.024.16 | 54.214.22 | 20.982.65 | 72.392.33 | 79.882.27 | 80.012.67 |
| DRIVE[50] | DSC() | 78.753.30 | 70.824.03 | 17.313.19 | 80.524.13 | 82.903.84 | 83.563.25 | |
| CHASEDB1[51] | DSC() | 70.965.53 | 68.672.19 | 23.302.54 | 74.841.91 | 81.152.21 | 81.522.06 | |
| EM | CREMI[14] | DSC() | 68.689.77 | 61.049.49 | 23.234.88 | 69.219.77 | 78.709.40 | 79.099.16 |
| Modality | Dataset | Metric | MAE | MFM | CLIP | MedSAM | FrepaViT | FrepaSwinT |
| CT | Pulmonary artery[15] | DSC() | 62.5913.22 | 57.9712.66 | 20.1913.55 | 64.5711.37 | 77.579.53 | 79.5810.69 |
| Pulmonary vein[15] | DSC() | 57.5710.37 | 52.4411.73 | 18.3911.22 | 61.7111.70 | 75.6510.99 | 77.852.94 | |
| Abdomen1K-Liver[45] | DSC() | 94.002.02 | 91.661.64 | 70.8710.48 | 94.881.38 | 95.141.37 | 94.541.96 | |
| Abdomen1K-Kidney[45] | DSC() | 89.958.69 | 79.1223.29 | 26.7817.54 | 91.185.17 | 92.154.87 | 92.473.87 | |
| Abdomen1K-Spleen[45] | DSC() | 91.805.61 | 89.077.08 | 43.8211.11 | 94.901.50 | 95.361.11 | 95.501.55 | |
| MRI | AMOS22[46] | DSC() | 71.235.13 | 70.606.27 | 26.0615.26 | 73.983.18 | 76.394.34 | 77.134.72 |
| ADCD[47] | DSC() | 68.2718.66 | 59.3622.50 | 18.371.07 | 71.7418.89 | 80.5419.57 | 82.9915.03 |
4.5 Ablation study
We gradually perform the ablation study of the components in Frepa to evaluate their respective contributions. The validation is performed on datasets including Pneumonia MNIST [12] for classification, Lung cancer [37] for lesion detection, and OCTA arteries and veins [48] for segmentation. The four most critical components of Frepa examined are frequency dual-component masking (FM), equalized-histogram image-domain masking (IM), HFL loss, and adversarial training (AT). For the ablation, these components are individually replaced or removed as follows: FM branch is deleted, IM is replaced by zero-patch masking (Fig. 3, second column), HFL loss, and AT are excluded during the pretraining phase.
| Pneumonia MNIST | Lung cancer | OCTA artery | OCTA vein | |||
| Metric | ACC | AUC | Detetcion | IoU | DSC() | DSC() |
| Frepa+ViT | 0.934 | 0.970 | 0.951 | 0.656 | 78.00 | 76.15 |
| w/o FM | 0.921 | 0.965 | 0.916 | 0.590 | 73.48 | 71.71 |
| w/o IM | 0.931 | 0.972 | 0.919 | 0.607 | 76.31 | 73.89 |
| w/o HFL | 0.927 | 0.967 | 0.939 | 0.624 | 74.69 | 73.11 |
| w/o AT | 0.933 | 0.970 | 0.945 | 0.632 | 77.70 | 76.03 |
| Frepa+SwinT | 0.932 | 0.981 | 0.948 | 0.677 | 79.08 | 77.35 |
| w/o FM | 0.913 | 0.971 | 0.905 | 0.604 | 73.96 | 72.40 |
| w/o IM | 0.926 | 0.977 | 0.919 | 0.621 | 76.58 | 76.65 |
| w/o HFL | 0.930 | 0.978 | 0.928 | 0.624 | 76.23 | 75.90 |
| w/o AT | 0.932 | 0.980 | 0.933 | 0.655 | 78.82 | 76.28 |
The results of the ablation study are presented in Table V. Removing each component leads to a decline in model performance across all tasks. Notably, the FM has the most significant impact; its removal causes the performance to degrade to be comparable to that of the MAE, suggesting a failure in high-frequency representation. Similarly, the results indicate that the zero-patch masking is less effective than our proposed equalized-histogram IM strategy. Additionally, the exclusion of HFL loss and AT also contributes to a reduction in performance. These findings are consistent for both ViT and SwinT architectures. In summary, the masking strategy and loss function designs are crucial components that significantly enhance the performance of Frepa.
4.6 Frequency representation in image embedding
In this section, instead of using downstream tasks, we design a more straightforward approach to evaluate the representation and preservation of high-frequency components in the image embeddings. The experimental pipeline is illustrated in Fig. 6a. In this setup, raw images are fed into fixed encoders (offline), and then decoders (online) are trained to reconstruct the raw images
MSE loss is employed during training. Subsequently, we compare the reconstructed images with the raw images across different frequency components, utilizing the RMSE as the metric for reconstructed ratios.
where represents low-pass, medium-pass, and high-pass filtering, respectively. We compared Frepa, implemented with both ViT and SwinT, against two other autoencoder methods: MAE and MFM. The results of this comparison are presented in Fig. 6b. It is evident that Frepa consistently outperforms the other autoencoder methods in medium-frequency and high-frequency components while maintaining comparable performance in low-frequency components. This underscores Frepa’s superior ability to represent and preserve high-frequency information without compromising low-frequency representation. These findings align with our experiments on downstream tasks, providing a direct and intuitive evaluation of the level of embedding spaces.
5 Discussion and Conclusion
Foundation models have achieved notable success with natural images by allowing the encoder to understand images from a global-level perspective. This approach is beneficial for tasks such as classification and object detection. However, in medical image analysis, this method faces significant challenges due to the anatomical similarities and low intra-modality variance among individuals [52], which result in less significance of global-level perception within one modality. Additionally, accurate diagnosis commonly depends on detailed information from specific anatomical structures, requiring more refined representations. Consequently, capturing only global-level and low-frequency information is inadequate for medical image analysis.
Our proposed Frepa addresses these limitations by enhancing high-frequency components and fine-grained features without sacrificing image-level perception. Our experiments demonstrate that Frepa can achieve a general performance improvement, especially for those requiring high-frequency details, surpassing both SOTA models and task-specific models. Our experiments further directly demonstrate that Frepa better captures and preserves high-frequency components from various perspectives, suggesting its potential use in other fields requiring fine-grained information, such as geospatial learning and spectral imaging. While our study shows excellent performance, future work could focus on optimizing Frepa for coarse-grained tasks and enhancing robustness to image noise, particularly in low-dose CT scans. Frepa’s versatile and effective approach paves the way for advancements in medical AI foundation models.